OpenStudy (anonymous):

HELPPPPPP Write an exponential function to represent the spread of Ben's social media post. Write an exponential function to represent the spread of Carter's social media post.

OpenStudy (anonymous):

Ben shared his post with 2 friends. Each of those friends shares with 3 more every day, so the number of shares triples every day. Carter shared his post with 10 friends, who each share with only 2 people each day.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@carolinar7

OpenStudy (anonymous):

@pooja195

OpenStudy (carolinar7):

@andrearchat

OpenStudy (carolinar7):

@laughoutloud

OpenStudy (laughoutloud):

I think the answer is on here http://www.peeranswer.com/study/mathematics/55ad36037590c3027a91bacc

OpenStudy (anonymous):

@campbell_st can u help me

OpenStudy (anonymous):

Ben shared his post with 2 friends. Each of those friends shares with 3 more every day, so the number of shares triples every day.

OpenStudy (campbell_st):

so the pattern is 2, 6, 18, 54, ..... initally ben has 2 friends f(x) = 2(a)^x a is the base and x is the time.... in days so if things triple, what number do you think will be the base..?

OpenStudy (campbell_st):

for carter, the pattern is 10, 20, 40, 80, ..... so g(x) = 10(b)^x b is the base and x is the time in days so what do you think the base is for this problem if its doubling..?

OpenStudy (anonymous):

I know carter is 10 (2)^x

OpenStudy (anonymous):

So for ben it would be 10(3)^x

OpenStudy (anonymous):

right @campbell_st

OpenStudy (campbell_st):

the base is 3... but its not 10 in front.... go and look at what I posted for Ben

OpenStudy (anonymous):

sry 2(3)^x

OpenStudy (campbell_st):

that would be my answer... when x = 0 which is the initial time, ben has 2 friends f(0) = 2*(3)^0 = 2